A metal alloy is 25% copper. Another metal alloy is 50% copper. How much of each should be used to make 1000 grams of a metal that is 45% copper? Show your system of equations and all work!
Answer by mananth(16949)

percent ---------------- quantity
Metal type I 25.00% ---------------- x grams
Metal type II 50.00% ------ 1000 - x grams
Mixture 45.00% ---------------- 1000
12
25.00% x + 50.00% ( 1000 - x ) = 45.00% * 1000
25 x + 50 ( 1000 - x ) = 45000
25 x + 50000 - 50 x = 45000
25 x - 50 x = 45000 - -50000
-25 x = -5000
/ -25
x = 200 grams 25.00% Metal type I
800 grams 50.00% Metal type II

m.ananth@hotmail.ca